OpenStudy (anonymous):

Solve the equation: p^2 =23 p -132

OpenStudy (anonymous):

GO SEE THE ANSWER

OpenStudy (anonymous):

go see the answer where?

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First rearrange it so it's equal to 0: \[p^2-23p+132=0\]Then use the quadratic equation to solve for p:\[p=\frac{-b \pm \sqrt{b^2-4ac}}{2a} = \frac{-(-23) \pm \sqrt{23^2-4(1)(132)}}{2(1)} = \frac{23}{2} \pm \frac{1}{2}\] So p is 11 and 12 Because these are integers, you know you could have solved it by factoring, without the quadratic equation. It can be factored as:\[p^2-23p+132 = (p-11)(p-12)\] 11 and 12 would make these 0, so they are your roots.

OpenStudy (anonymous):

thank you so much